hi!
my problem is now: THE DISTANCE BETWEEN THE ROAD OF A BRIDGE AND THE TOP PART OF ITS ARCH IS 3m. IF THE ARCH OF THE BRIDGE IS GIVEN BY THE EQUATION
h(x)=7x-x^2, FIND THE HEIGHT OF THE BRIDGE.
thank you!

Answer by stanbon(75887)

THE DISTANCE BETWEEN THE ROAD OF A BRIDGE AND THE TOP PART OF ITS ARCH IS 3m. IF THE ARCH OF THE BRIDGE IS GIVEN BY THE EQUATION
h(x)=7x-x^2, FIND THE HEIGHT OF THE BRIDGE.
----------------
Maximum height of the arch occurs at x = -b/2a = -7/(2*-1) = (7/2)
The height of the arch is f(7/2) = 7(7/2)-(7/2)^2 = 49/2 - 49/4 = 49/4= 12.25m
-------------------
The height of the bridge is 12.25-3 = 9.25 meters
======================
Cheers,
Stan H.